What is the magnitude of a point charge chosen so that the electric field 50 cm away has magnitude 2N/C ?

Grade:11

1 Answers

Arun
6 years ago
k = 9 * 10 ^ 9 
r = 0.5 m 
E = K q/r^2 = 2 N/c 
q = E * r ^ 2 / k 
= 2 * 0.25 / 9* 10^9 C
 = 0.5 * 10^-9 /9
= 5.5 * 10^-11 c